OpenStudy (anonymous):

For the function f(x) = |2x - 7| + 3, find f(-2)

OpenStudy (anonymous):

plug in -2 for all x values

OpenStudy (anonymous):

Let's see... f(-2) = |2(-2) - 7| + 3 f(-2) = |-4 - 7| + 3 f(-2) = |-11| + 3 f(-2) = 11 + 3 f(-2) = 14

OpenStudy (anonymous):